If fixed costs are $282,000, the unit selling price is $31, and the unit variable costs are $18, what are the break-even sales (units) if fixed costs are reduced by $33,300?

a.15,305 units

b.22,957 units

c.28,696 units

d.19,131 units
Answer #1

New fixed costs=282,000-33,300=$248700

Contribution margin=Sales-Variable cost

=31-18=$13 per unit

Breakeven=fixed costs/Contribution margin

=$248700/13

=19131 units(Approx).